OpenStudy (technodynamic):

Let o(x)=sqrt(3x+8) . Find a simplified formula for the function below. Your final answer should be a polynomial with only one term in each power of x. (o(x))^2. Can someone explain and show step-by-step how to do this?

OpenStudy (anonymous):

o(x)=sqrt(3x+8) ( o(x) )^2 = ( sqrt(3x+8) )^2 ( o(x) )^2 = 3x + 8